### v2

# smalltableaux

Abstract
Kostka, Littlewood-Richardson, Plethysm and Kronecker coefficients are multiplicities of irreducible representations (irreps) of the symmetric group in restrictions and products of irreps. They play an important role in representation theory and are notoriously hard to compute. We give quantum algorithms that efficiently compute these coefficients whenever the ratio of dimensions of the representations is polynomial. Using that the Kostka numbers admit combinatorial interpretation, we show that there is an efficient classical algorithm for polynomially-bounded Kostka numbers and conjecture existence of a similar algorithm for the Littlewood-Richardson coefficients. We argue why the same classical algorithm does not straightforwardly work for the Plethysm and Kronecker coefficients, give evidence on how our quantum algorithm may avoid some hardness obstructions in their computation, and conjecture that the problem could lead to superpolynomial quantum speedups on some inputs. We finally use Frobenius reciprocity to derive another quantum algorithm that estimates these coefficients using induction and has a different cost-to-input dependence.
Table 1: Quantum Algorithms for Branching Coefficients. Choices of groups $H\subseteq G$, $H$-representation $R$, and an $H$-irrep $r^{\alpha}_{H}$ for which the algorithm in Thm. 1 computes the studied multiplicity coefficients. Here, $n,a,b,c,d\in\mathbb{N}$ where $a+b=n$ (Littlewood-Richardson) and $cd=n$ (Plethysm). The cost $\dim(R)/\dim(r^{\alpha}_{H})$ is the number of samples required to (exactly) compute the coefficient .
Problem Input
Group
$G$
Subgroup
$H\subseteq G$
$H$-Rep
$R$
$H$-Irrep
$r^{\alpha}_{H}$
Output
${\rm\mathbb{mult}}(r^{\alpha}_{H},R)$
Cost
$\mathcal{O}(\frac{\dim(R)}{\dim(r^{\alpha}_{H})})$
Kostka $\nu,\mu\vdash n$ $S_{n}$ $S_{\mu}:=\bigtimes_{i}S_{\mu_{i}}$ $(r^{\nu}_{S_{n}})\downarrow^{S_{n}}_{S_{\mu}}$ $\bigotimes_{i}r_{S_{\mu_{i}}}^{(\mu_{i})}$ $K_{\nu}^{\mu}$ $\mathcal{O}\big{(}d_{\nu}\big{)}$
Littlewood $\nu\vdash n$ and $\lambda,\mu\vdash a,b$ $S_{n}$ $S_{a}\times S_{b}$ $(r_{S_{n}}^{\nu})\downarrow^{S_{n}}_{S_{a}\times S_{b}}$ $r^{\lambda}_{S_{a}}\otimes r^{\mu}_{S_{b}}$ $c_{\lambda\mu}^{\nu}$ $\mathcal{O}\big{(}\frac{d_{\nu}}{d_{\lambda}d_{\mu}}\big{)}$
Plethysm $\nu\vdash n$ and $\lambda,\mu\vdash c,d$ $S_{n}$ $S_{c}\wr S_{d}$ $(r^{\nu}_{S_{n}})\downarrow^{S_{n}}_{S_{c}\wr S_{d}}$ $r^{\lambda}_{S_{c}}\wr r^{\mu}_{S_{d}}$ $a_{\lambda\mu}^{\nu}$ $\mathcal{O}\big{(}\frac{d_{\nu}}{d_{\lambda}^{b}d_{\mu}}\big{)}$
Kronecker [1] $\nu,\lambda,\mu\vdash n$ $S_{n}\times S_{n}$ $S_{n}$ $(r^{\lambda}_{S_{n}}\otimes r^{\mu}_{S_{n}})\downarrow^{S_{n}\times S_{n}}_{S_{n}}$ $r_{S_{n}}^{\nu}$ $g_{\lambda\mu\nu}$ $\mathcal{O}\big{(}\frac{d_{\lambda}d_{\mu}}{d_{\nu}}\big{)}$
Introduction.–
[…]
Let $G$ be a finite group. A unitary representation $R$ of $G$ on a finite-dimensional vector space $\mathcal{H}=\mathbb{C}^{d}$ is a function $R$ from $G$ to the group of unitary matrices on $\mathcal{H}$ such that $R(g)R(h)=R(gh)$. Maschke’s theorem asserts that for any $G$-representation $R$ there exists a unitary $U$ that simulaneously block-diagonalizes all $R(g)$ as:
$UR(g)U^{\dagger}=\bigoplus_{\alpha}I_{{\rm\mathbb{mult}}(r^{\alpha}_{G},R)}\otimes r^{\alpha}_{G}(g),\,$ (1)
[…]
We give quantum algorithms for computing the following multiplicity coefficients:
  1. 1.
    Kostka Numbers, the multiplicity of the trivial $S_{\mu}$-irrep in the restriction of an $S_{n}$-irrep:
    $K^{\mu}_{\nu}={\rm\mathbb{mult}}(r^{\rm triv}_{S_{\mu}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{\mu}}).$ (2)
    where $\nu,\mu\vdash n$ and $S_{\mu}=\bigtimes_{i}S_{\mu_{i}}$ is a Young subgroup. $K^{\mu}_{\nu}$ also counts the number of semistandard Young tableaux of shape $\nu$ and content $\mu$. They play role in quantum tomography (see for example Ref. [6]).
  2. 2.
    Littlewood-Richardson Coefficients, the multiplicty of $S_{a}\times S_{b}$-irreps in an $S_{n}$-irrep for $a+b=n$:
    $c^{\nu}_{\lambda\mu}$ $={\rm\mathbb{mult}}(r^{\lambda}_{S_{a}}\otimes r^{\mu}_{S_{b}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{a}\times S_{b}}).$ (3)
    By Schur-Weyl duality [7, 8], this is also a multiplicty of $\textsf{SU}(d)$-irreps in the tensor product of two other $\textsf{SU}(d)$-irreps. The coefficient plays a role in angular momentum recoupling theory (for $d=2$, see for example Refs. [9, 10, 11]) or particle physics (for example for $d=3$, the “eight-fold” way [12]).
  3. 3.
    Plethysm Coefficients, the multiplicity of irreps of the wreath product subgroup $S_{c}\wr S_{d}\subset S_{n}$ for $cd=n$ in an $S_{n}$-irrep:
    $a^{\nu}_{\lambda\mu}$ $={\rm\mathbb{mult}}(r^{\lambda\mu}_{S_{c}\wr S_{d}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{c}\wr S_{d}})\,,$ (4)
    where $\lambda,\mu,\nu\vdash c,d,n$. It is also the multiplicity of $\textsf{SU}(d)$-irreps in the composition of a pair of them. They appear in the quantum marginal problem for distinguishable particles [13, Sec. 7].
Our construction also subsumes an algorithm for Kronecker coefficients which previously appeared in Ref. [1]. Kronecker coefficients are multiplicities of $S_{n}$-irreps in the tensor products of $S_{n}$-irreps:
$g_{\lambda\mu\nu}$ $={\rm\mathbb{mult}}(r^{\nu}_{S_{n}},r^{\lambda}_{S_{n}}\otimes r^{\mu}_{S_{n}}),$ (5)
and play role in the quantum marginal problem [8, 14].
Quantum Algorithms.–
Computation of the above coefficients is worst-case ##\#$\P$# ¶-hard [15, 16, 17] and we thus do not expect to find a quantum algorithm that is efficient on all inputs. In what follows, we give a quantum algorithm that is efficient on some inputs and analyze its performance.
Let $G$ be a finite group and $R$ be a $G$-representation on $\mathbb{C}[G]$ with orthonormal basis {|$\{\ket{g}\}_{g\in G}$ acting as:
$R(h)\ket{g}$ =|$=\ket{hg},\quad\text{for all $h,g\in G$.}$ (6)
Such representation is called left-regular representation. A transformation $U$ that maximally block-diagonalizes this representation as in Eq. (1) is the Quantum Fourier Transform (QFT); we write $G$-QFT whenever we need to make the dependence on $G$ explicit. It is defined as:
$\langle\alpha,i,j|U|g\rangle$ $=\sqrt{\frac{d_{\alpha}}{|G|}}\innerproduct{i}{r_{G}^{\alpha}(g)|j},$ (7)
where $r_{G}^{\alpha}(g)$ is a $G$-irrep labelled by $\alpha$ and $\innerproduct{i}{r_{G}^{\alpha}(g)|j}$ for $i,j\in[d_{\alpha}]$ is its $i,j$-th matrix element. Efficient quantum algorithms for $G$-QFT are known for $G=S_{n}$ [18] and some other finite groups  [19]. We use QFT over the Young subgroup $S_{\mu}=\bigtimes_{i}S_{\mu_{i}}$ and the wreath product group $S_{a}\wr S_{b}$ from Ref. [19]. See Supp. Mat. for details.
Our algorithm relies on the efficient preparation of states of the form:
$\rho^{\alpha}_{G}$ $=U^{\dagger}\left(\ket{\alpha}\bra{\alpha}\otimes\frac{I_{d_{\alpha}}}{d_{\alpha}}\otimes\frac{I_{d_{\alpha}}}{d_{\alpha}}\right)U=\frac{\Pi^{G}_{\alpha}}{d_{\alpha}^{2}},$ (8)
where $I_{d_{\alpha}}$ is the $d_{\alpha}\times d_{\alpha}$ identity matrix and $\Pi^{G}_{\alpha}$ is the projector onto the maximally isotypic subspace $\alpha$ in the regular representation of $G$. The maximally isotypic subspace of a given type is the subspace of $\mathcal{H}$ containing all irreps of such type. We note that an efficient $G$-QFT is a sufficient (but possibly not necessary) to prepare such a state. Equipped with this, we show:
Theorem 1.
Let $H\subseteq G$. If there is an efficient quantum circuit for $H$-QFT, and an efficient way of preparing $\rho^{\alpha}_{G}$ in Eq. (8), then there is a quantum algorithm that computes ${\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow_{H}^{G})$ in time $poly(n,d_{\alpha}/d_{\beta})$.
Proof.
See Fig. 1. The algorithm is a variation on Harrow’s generalized phase estimation [20]. It uses two registers: a target register $\mathcal{T}$ and a control register $\mathcal{C}$. It inputs $\rho^{\alpha}_{G}$ from Eq. (8) to the target register and a uniform superposition over the subgroup $H\subseteq G$ to the control register. It then applies a controlled action of the $G$-regular representation $R$ to $\rho^{\alpha}_{G}$. Since $R$ is controlled by a register that contains a superposition over the subgroup $H$, it acts as a restricted representation $R\downarrow_{H}^{G}$ which decomposes into irreducible representations of $H$. The algorithm concludes by measuring the $H$-irrep label on the control register. This is described by:
$E_{\beta}=\frac{1}{\sqrt{|H|}}\sum_{h\in H}\Bigg{[}(\outerproduct{\beta}{\beta}\otimes I_{d_{\beta}}\otimes I_{d_{\beta}})V\ket{h}\Bigg{]}_{\mathcal{C}}\otimes\Bigg{[}R(h)\Bigg{]}_{\mathcal{T}}$ (9)
where $V$ is the $H$-QFT, $\Pi^{H}_{\beta}$ is the projector on the $\beta$-isotypic subspace in the regular representation of $H$, $\alpha$ labels irreps of $S_{n}$. The measurement outcome $\beta$ appears with probability:
$p_{\alpha}(\beta)$ $=\text{Tr}\left[E_{\beta}\rho^{\alpha}_{S_{n}}E_{\beta}^{\dagger}\right]={\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow^{G}_{H})\frac{d_{\beta}}{d_{\alpha}}.$ (10)
Because ${\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow^{G}_{H})$ is an integer, we can get its exact value in ${\rm poly}(d_{\beta}/d_{\alpha})$ many shots with high probability by the Chernoff-Hoeffding bound. It follows that the algorithm computes ${\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow^{G}_{H})$ in ${\rm poly}(n,d_{\alpha}/d_{\beta})$ time. See the Supp. Mat. for more details. ∎
Tab. 1 summarizes different choices of $H$ and $G$ where the quantum algorithm in Thm. 1 computes the coefficients in Eqs. (2) to (5), which we now detail.
Figure 1: The quantum algorithm in Thm. 1. The algorithm initializes a maximally-mixed state over the isotypic subspace $\alpha$ in the $G$-regular representation, applies a $G$-regular representation controlled by a uniform superposition over $H$ and weak-Fourier samples the control register.
Kostka Numbers.—
The quantum algorithm for the Kostka numbers fixes $H=S_{\mu}\subseteq G=S_{n}$ for a partition $\mu\vdash n$ where $S_{\mu}$ is the Young’s subgroup $S_{\mu}:=\bigtimes_{i}S_{\mu_{i}}$. The Kostka number $K_{\lambda}^{\mu}$ is the multiplicity of the trivial irreducible representation in the restriction of an $S_{n}$-irrep $r^{\nu}_{S_{n}}$ to $S_{\mu}$, $K^{\mu}_{\nu}={\rm\mathbb{mult}}(r_{S_{\mu}}^{{\rm triv}},r^{\nu}_{S_{n}}\downarrow_{S_{\mu}}^{S_{n}})$.
An efficient $S_{\mu}$-QFT can be setup from a tensor product of $S_{\mu_{i}}$-QFTs for each factor in $\mu\vdash n$. The probability distribution sampled by the algorithm (Eq. (10)) gives the label of the trivial irreducible representation of $S_{\mu}$ with probability:
$p_{\nu,\mu}({\rm triv})$ $=K^{\mu}_{\nu}/d_{\nu},$ (11)
which suffices to compute the Kostka number efficiently whenever $d_{\nu}\leq{\rm poly}(n)$.
Computing Kostka numbers is a ##\#$\P$# ¶-complete problem [15, 21], but one may hope that restricting the problem to $K^{\mu}_{\nu}$ for $d_{\nu}\leq{\rm poly}(n)$, where Thm. 1 gives an efficient quantum algorithm, remains hard for classical algorithms. We now show that this is not the case.
A semistandard Young tableau $T$ of shape $\nu$ and content $\mu$ is a tableau where an integer $i$ appears $\mu_{i}$-many times such that the entries are nondecreasing along each row and strictly increasing along every column. The Kostka number $K^{\mu}_{\nu}$ counts the number of semistandard Young tableaux of shape $\nu$ and content $\mu$. It is known that $\lambda\succeq\mu\implies K_{\nu}^{\lambda}\geq K_{\nu}^{\mu}$, where $\succeq$ is the dominance order (meaning $\nu_{1}+\ldots+\nu_{r}\geq\mu_{1}+\ldots+\mu_{r}$ for all $r$) [22]. This particularly means that $d_{\nu}\geq K_{\nu}^{\mu}$ for all $\mu$, which implies that in the regime where the quantum algorithm is efficient, $K^{\mu}_{\nu}$ counts only a polynomial number of tableaux. We use this to show that:
Theorem 2.
There is a deterministic polynomial time algorithm for computing the Kostka number $K^{\mu}_{\nu}$ for $\mu,\nu\vdash n$ such that $d_{\nu}\leq{\rm poly}(n)$.
The proof is a polynomial-time classical algorithm that uses dynamic programming. See the Supp. Mat.
Littlewood-Richardson Coefficients.–
The quantum algorithm for Littlewood-Richardson (LR) coefficients uses the same circuit as the algorithm for the Kostka numbers but fixes $H=S_{a}\times S_{b}\subseteq G=S_{n}$ for $a+b=n$. The irreducible representations of $S_{a}\times S_{b}$ are tensor product of irreducibles $r_{S_{a}}^{\lambda}\otimes r_{S_{b}}^{\mu}$, labelled by pairs $\alpha=(\lambda,\mu)$ with $\lambda,\mu\vdash a,b$. We have $c^{\nu}_{\lambda\mu}={\rm\mathbb{mult}}(r^{\lambda}_{S_{a}}\otimes r^{\mu}_{S_{b}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{a}\times S_{b}})$, and from Thm. 1, the algorithm samples from:
$p_{\nu}(\lambda,\mu)$ $=c^{\nu}_{\lambda\mu}\frac{d_{\lambda}d_{\mu}}{d_{\nu}}.$ (12)
This suffices to compute the LR coefficient whenever $d_{\nu}/d_{\lambda}d_{\mu}\leq{\rm poly}(n)$.
LR coefficients count combinatorial objects called the LR tableaux. A skew shape $\nu/\lambda$ is a diagram where $\lambda$ is substracted from $\nu$ and a skew semistandard Young Tableau is a skew shape filled with integers such that the entries are non-decreasing along every row and strictly increasing along every column. The LR tableau is a skew Young tableau of shape $\nu/\lambda$ whose row-reading evaluates to $\mu$. The LR coefficient $c^{\nu}_{\lambda\mu}$ is the number of such tableaux; this fact is called the LR rule. See [23, 24, 25] and the Supp. Mat. for details (esp. definition of ‘row reading’). The quantum algorithm in Thm. 1 is only efficient for $c^{\nu}_{\lambda\mu}\leq{\rm poly}(n)$ in which case it counts only up to ${\rm poly}(n)$ many LR tableaux. In analogy to Thm. 2, we expect that there is a classical algorithm that efficiently computes LR coefficients under this constraint:
Hypothesis 1.
There is a polynomial-time algorithm for computing the Littlewood-Richardson coefficient $c^{\nu}_{\lambda\mu}$ for $\nu\vdash n$, $\lambda\vdash a$ and $\mu\vdash b$ for $a+b=n$ and $d_{\nu}/d_{\lambda}d_{\mu}\leq{\rm poly}(n)$.
If true, this rules out superpolynomial speedups for computing LR coefficients by Thm. 1.
Plethysm Coefficients.— The algorithm for Plethysm coefficients sets $H=S_{c}\wr S_{d}\subseteq G=S_{n}$ for $cd=n$. An efficient implementation of the QFT over this group is known only for $|S_{a}|\leq{\rm poly}(n)$ [19]. Under this restriction, the quantum algorithm in Thm. 1 samples efficiently from a distribution:
$p_{\nu}(\lambda,\mu)$ $=a^{\nu}_{\lambda\mu}\frac{d_{\lambda}d_{\mu}^{d}}{d_{\nu}},$ (13)
where $a^{\nu}_{\lambda\mu}={\rm\mathbb{mult}}(r^{\lambda,\mu}_{S_{c}\wr S_{d}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{c}\wr S_{d}})$ is the Plethysm coefficient.
It is not known if Plethysm coefficients count a well defined set of combinatorial objects (Prob. 9 on Stanley’s list [26]), which means that the simulation strategy of Thm. 2 doesn’t straightforwardly apply [27, 17, 28]. The classical complexity of computing the Plethysm coefficients was recently studied in Ref. [17] where it was shown that the problem is in $\P$ if the number of parts $\ell(\lambda)$ is constant and $\mu={\rm triv}$, but becomes ##\#$\P$# ¶-hard even for fixed $d=3$ and $\ell(\lambda)$ unrestricted. While this suggests that the complexity of the problem is controlled by the parameter $c$, its complexity for $|S_{c}|={\rm poly}(n)$ ($c\sim\log(n)$) remains to our best knowledge open. We thus state:
Conjecture 1.
There is no polynomial time randomized algorithm that (with high probability) computes $a^{\nu}_{\lambda\mu}$ for $\lambda\vdash c$ and $\mu\vdash d$, such that $cd=n$, $c\leq{\rm poly}(\log(n))$ and $d_{\lambda}d_{\mu}^{d}/d_{\nu}\geq 1/{\rm poly}(n)$.
If this conjecture was true, the quantum algorithm for Plethysm coefficient gives a superpolynomial quantum speedup. The fact that an analog of Thm. 2 does not apply here gives some support for this, but the constraints on the conjectured computationally hard regime for the problem are very stringent and can make the restricted computational problem significantly easier. Beyond that, we have little intuition whether the conjecture is true or false.
Kronecker Coefficients.— Given $\lambda,\mu,\nu\vdash n$, the Kronecker coefficient $g_{\lambda\mu\nu}$ is defined as $g_{\lambda\mu\nu}={\rm\mathbb{mult}}(r^{\lambda}_{S_{n}}\otimes r^{\mu}_{S_{n}},r^{\nu}_{S_{n}})$. The algorithm in Ref. [29] is a special case of Thm. 1, for the choice $H=S_{n}\subseteq G=S_{n}\times S_{n}$. It samples from the output distribution:
$p_{\lambda\mu}(\nu)$ $=\frac{d_{\nu}}{d_{\lambda}d_{\mu}}g_{\lambda\mu\nu}\,,$ (14)
and computes $g_{\lambda\mu\nu}$ whenever $d_{\lambda}d_{\mu}/d_{\nu}\leq{\rm poly}(n)$ or (a weaker condition) $d_{\nu}\leq{\rm poly}(n)$.
Similarly to Plethysms, it remains a major open question if the Kronecker coefficients count some simple set of combinatorial objects (Stanley’s Prob. 10 [26]), which again precludes straightforward application of the simulation strategy of Thm. 2. The classical complexity of computing the Kronecker coefficients was studied in Ref. [28] which identified the extensive number of parts as a potential obstruction to its efficient computation. They show that $g_{\lambda\mu\nu}$ can be computed in time:
$\mathcal{O}\Big{(}L\log(N)\Big{)}+\Big{(}L\log(M)\Big{)}^{\mathcal{O}(L^{3}\log L)},$ (15)
where $L=\max\Big{(}\ell(\lambda),\ell(\mu),\ell(\nu)\Big{)}$, $N=\max(\lambda_{1},\mu_{1},\nu_{1})$ is the largest length of the first row and $M=\max(\lambda_{2},\mu_{2},\nu_{2})$ is the largest length of the second row among these partitions [28, Thm. 3.5]. Even for fixed $L$, the algorithms of Refs. [30, 31] compute the Kronecker coefficient in polynomial time but become inefficient when the maximum length of the input partitions varies with input size. The algorithm in Ref. [32] computes Kronecker coefficient by reduction to vector partition function problem. All of these algorithms are based on Barvinok’s algorithm for counting lattice points in convex polytopes (Refs. [33, 34, 35, 36, 37, 38, 39]). The quantum algorithm in Thm. 1 is, for all practical purposes, outperformed by such algorithms on some inputs. For example, Ref. [32, Ex. 3.4] computes $g_{\lambda\mu\nu}=391$ for $\lambda=[140^{2}],\;\mu=[70^{4}],\;\nu=[57^{3},33^{3},10]$ for which $d_{\nu}/(d_{\lambda}d_{\mu})\approx 1.8\times 10^{-31}$. The quantum algorithm in Thm 1 requires an order of $10^{31}$ samples on this input, which makes it useless on this input. The length obstruction in Eq. (15) nonetheless supports the conjecture of Ref. [1] that the algorithm for the Kronecker coefficients can outperform classical algorithms for triples of partitions with large number of parts. We (re)state:
Conjecture 2 (See also [1]).
There is no polynomial time randomized algorithm that (with high probability) computes $g_{\lambda\mu\nu}$ for $\mu,\nu,\lambda\vdash n$ and $d_{\nu}\leq{\rm poly}(n)$.
If true, the quantum algorithm of Thm. 1 gives a superpolynomial quantum speedup. As long as $\lambda,\mu$ have number of parts that scale with $n$, the quantum algorithm can do better than the bound in Eq. (15) as well as the state of the art classical algorithms. This of course assumes that there exist partitions with extensive number of parts that satisfy $d_{\nu}\leq{\rm poly}(n)$ on which the computation of $g_{\mu\nu\lambda}$ remains nontrivial; refuting their existence would also imply that that the Conj. 2 is false. See the Supp. Mat. for a numerical study and additional discussion.
Sylow Branching.–
We remark that Thm. 1 possibly extends to other interesting coefficients. For example, the works of [40, 41] study restrictions of the $S_{n}$-irreducibles to Sylow $p$-subgroups. A Sylow $p$-subgroup is a subgroup of order $mp^{k}$ for some integer $m$ and the largest possible $k$. The Sylow branching coefficients are $z^{\nu}_{\alpha}={\rm\mathbb{mult}}(r^{\alpha}_{P_{n}},r^{\nu}_{S_{n}}\downarrow_{P_{n}}^{S_{n}}),$ where $\alpha$ labels $P_{n}$-irreps. If we could implement a QFT over a Sylow subgroup, Thm. 1 computes $z^{\nu}_{\alpha}$ whenever $d_{\alpha}/d_{\nu}\geq 1/{\rm poly}(n)$.
Induction Algorithm.–
Frobenius’s reciprocity [7]:
${\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow_{H}^{G})={\rm\mathbb{mult}}(r^{\alpha}_{G},r^{\beta}_{H}\uparrow_{H}^{G})\,,$ (16)
suggests an alternative to Thm. 1 that computes multiplicities by Fourier sampling over induced representations:
Theorem 3.
Let $H\subseteq G$. If there is an efficient quantum circuit for $G$-QFT and an efficient way of preparing $\rho^{\beta}_{H}$ in Eq. (8), then there is an efficient quantum algorithm that computes ${\rm\mathbb{mult}}(r^{\alpha}_{G},r^{\beta}_{H}\uparrow_{H}^{G})$ in time $poly(n,d_{\beta}|G|/d_{\alpha}|H|)$.
Proof.
The algorithm starts with two registers, an $H$-regular rep register $\mathcal{S}=\mathbb{C}[H]$ and a coset register $\mathcal{V}=\mathbb{C}[G/H]$. It prepares $\rho^{\beta}_{H}$ from Eq. (8) in $\mathcal{S}$ and applies the ‘embedding’ unitary $U_{E}$ to $\mathcal{S}\otimes\mathcal{V}$: $U_{E}\Big{(}\ket{t}_{\mathcal{V}}\otimes\ket{h}_{\mathcal{S}}\Big{)}=\ket{g},$ where $t$ is a representative of the coset of $H$ in $G$ and $g=th$ is an element of $G$. This unitary can be implemented efficiently as long as the multiplication of the group elements in $G$ is efficient. The algorithm applies a $G$-QFT to $\mathcal{S}\otimes\mathcal{V}$ and measures the irrep label. The outcome $\alpha$ appears with probability:
$p_{\beta}(\alpha)$ $=\frac{|H|d_{\alpha}}{|G|d_{\beta}}{\rm\mathbb{mult}}(r^{\alpha}_{G},r^{\beta}_{H}\uparrow_{H}^{G}).$ (17)
Following the analysis in Thm. 1, we can find ${\rm\mathbb{mult}}(r^{\alpha}_{G},r^{\beta}_{H}\uparrow_{H}^{G})$ in $poly(n,d_{\beta}|G|/d_{\alpha}|H|)$ with high probability. See Supp. Mat. for details. ∎
Figure 2: The quantum algorithm in Thm. 3. The algorithm applies the embedding unitary $U_{E}$ to the subgroup and coset register and weak Fourier samples.
The sampling cost of the restriction algorithm in Thm. 1 is $d_{\alpha}/d_{\beta}$, compared to $(d_{\beta}|G|)/(d_{\alpha}|H|)$ in Thm. 3 and the restriction is preferable whenever $(d_{\alpha}/d_{\beta})^{2}<|G|/|H|$. For $\beta={\rm triv}$, Eq. (17) is the distribution of Weak Fourier sampling from the hidden subgroup state, appearing in the context of the Hidden Subgroup Problem (HSP). The computational task in HSP is of course different from computing the multiplicity, but the connection suggests generalizations of HSP to $\beta\neq{\rm triv}$ we explore in future work. See Supp. Mat. for details.
Discussion.–
We studied quantum algorithms for Kostka, Littlewood-Richardson, Kronecker and Plethysm coefficients which are efficient on a certain subset of inputs. We gave a polynomia-time classical algorithm for computing of the Kostka numbers $K^{\mu}_{\nu}$ whenever $d_{\nu}\leq{\rm poly}(n)$. This rules out superpolynomial speedups for computing the Kostka numbers by the quantum algorithm in Thm 1. The key enabling feature was the combinatorial interpretation of the Kostka numbers and since Littlewood-Richardson coefficients also admit such interpretation, we conjectured the existence of a similar classical algorithm for them. We explained why this strategy does not straightforwardly extend to the Plethysm and Kronecker coefficients and conjectured that our quantum algorithms efficiently solves problems that are hard for randomized algorithms.
Going against this thesis is a result in Ref. [42] in which Jordan presented a quantum algorithm for computation of $S_{n}$ characters to a similar precision as our Thm. 1 and subsequently an efficient (and fairly non-trivial) classical algorithm based on Roichman rule [43]. It is quite possible that Conjectures 1, 2 will have similar fate. In the other direction, it may be worth trying to promote the ‘hardness obstructions’ in Refs. [28, 17] to cryptographic hardness assumptions, for example by formulating them as distributional decision problems and proving worst-average case hardness reductions, similarly to what Ref. [44] did for the discrete logarithm problem.
Recently, Ref. [45] showed that some of the multiplicity coefficients considered here appear naturally in spin chains. It remains an interesting question if this could provide more insight into this class of quantum algorithms.
Acknowledgments.–
We would like to thank David Gosset, Sergey Bravyi, Louis Schatzki, Marco Cerezo, Marcos Crichigno, Samuel Slezak and Diego Garcia-Martín for discussions. M.L. was supported by the Center for Nonlinear Studies at LANL and by the Laboratory Directed Research and Development (LDRD) program of LANL under project number 20230049DR.