### v4

# Quantum Algorithms for Representation-Theoretic Multiplicities

Abstract
Kostka, Littlewood-Richardson, Plethysm and Kronecker coefficients are the multiplicities of irreducible representations in the decomposition of representations of the symmetric group that play an important role in representation theory, geometric complexity and algebraic combinatorics. We give quantum algorithms for computing these coefficients whenever the ratio of dimensions of the representations is polynomial and study the computational complexity of this problem. We show that there is an efficient classical algorithm for computing the Kostka numbers under this restriction and conjecture the existence of an analogous algorithm for the Littlewood-Richardson coefficients. We argue why such classical algorithm does not straightforwardly work for the Plethysm and Kronecker coefficients and conjecture that our quantum algorithms lead to superpolynomial speedups for these problems. The conjecture about Kronecker coefficients was concurrently disproved by Greta Panova in [arXiv:2502.20253] with a classical solution which, if optimal, points to a $\mathcal{O}(n^{4+2k})$ vs $\tilde{\Omega}(n^{4k^{2}+1})$ polynomial gap in quantum vs classcial computational complexity for an integer parameter $k$.
Table 1: Quantum Algorithms for Branching Coefficients. Choices of groups $H\subseteq G$, $H$-representation $R$, and an $H$-irrep $r^{\alpha}_{H}$ for which the algorithm in Thm. 1 computes the studied multiplicity coefficients. Here, $n,a,b,c,d\in\mathbb{N}$ where $a+b=n$ (Littlewood-Richardson) and $cd=n$ (Plethysm). The cost $\dim(R)/\dim(r^{\alpha}_{H})$ is the number of samples required to (exactly) compute the coefficient.
Problem Input
Group
$G$
Subgroup
$H\subseteq G$
$H$-Rep
$R$
$H$-Irrep
$r^{\alpha}_{H}$
Output
${\rm\mathbb{mult}}(r^{\alpha}_{H},R)$
Cost
$\mathcal{O}(\frac{\dim(R)}{\dim(r^{\alpha}_{H})})$
Kostka $\nu,\mu\vdash n$ $S_{n}$ $S_{\mu}:=\bigtimes_{i}S_{\mu_{i}}$ $(r^{\nu}_{S_{n}})\downarrow^{S_{n}}_{S_{\mu}}$ $\bigotimes_{i}r_{S_{\mu_{i}}}^{(\mu_{i})}$ $K_{\nu}^{\mu}$ $\mathcal{O}\big{(}d_{\nu}\big{)}$
Littlewood $\nu\vdash n$ and $\lambda,\mu\vdash a,b$ $S_{n}$ $S_{a}\times S_{b}$ $(r_{S_{n}}^{\nu})\downarrow^{S_{n}}_{S_{a}\times S_{b}}$ $r^{\lambda}_{S_{a}}\otimes r^{\mu}_{S_{b}}$ $c_{\lambda\mu}^{\nu}$ $\mathcal{O}\big{(}\frac{d_{\nu}}{d_{\lambda}d_{\mu}}\big{)}$
Plethysm $\nu\vdash n$ and $\lambda,\mu\vdash c,d$ $S_{n}$ $S_{c}\wr S_{d}$ $(r^{\nu}_{S_{n}})\downarrow^{S_{n}}_{S_{c}\wr S_{d}}$ $r^{\lambda}_{S_{c}}\wr r^{\mu}_{S_{d}}$ $a_{\lambda\mu}^{\nu}$ $\mathcal{O}\big{(}\frac{d_{\nu}}{d_{\lambda}^{d}d_{\mu}}\big{)}$
Kronecker [1] $\nu,\lambda,\mu\vdash n$ $S_{n}\times S_{n}$ $S_{n}$ $(r^{\lambda}_{S_{n}}\otimes r^{\mu}_{S_{n}})\downarrow^{S_{n}\times S_{n}}_{S_{n}}$ $r_{S_{n}}^{\nu}$ $g_{\lambda\mu\nu}$ $\mathcal{O}\big{(}\frac{d_{\lambda}d_{\mu}}{d_{\nu}}\big{)}$
Introduction.–
Is there a computational problem that admits efficient quantum algorithms but no efficient randomized one? Quantum computing is driven by the hypothesis that the answer is yes, to a large extent because – “…our best description of the physical world is quantum-mechanical and quantum mechanics seems hard to simulate by classical computers…” [2]. This intuition often comes with limited backing for specific computational problems and supporting it rigorously faces significant challenges in computational complexity theory. It would be ideal to prove that a decision problem admits efficient quantum algorithms but no randomized one, but doing that unconditionally entails showing that ${\mathsf{BPP}}\subset{\mathsf{BQP}}$ and in consequence separating ${\mathsf{BPP}}$ and ${\mathsf{PSPACE}}$, a major open problem in computational complexity theory. Here we resort to the next best thing: find a problem that could be outside ${\mathsf{BPP}}$; find an efficient quantum algorithm for it and study what makes it intractable classically. An excellent execution of this three-step process is Shor’s algorithm for factoring and discrete logarithm [3]: problems that are not known to be either ${\mathsf{NP}}$-hard or ${\mathsf{BQP}}$-complete, but the evidence for their hardness is that we use factoring and discrete logarithm for public-key cryptography [4, 5]. Here we give quantum algorithms for a class of problems in representation theory that could yield computational speedups, if our hardness conjectures could be similarly promoted to cryptographic hardness assumptions.
Let $G$ be a finite group. A unitary representation $R$ of $G$ on a finite-dimensional vector space $\mathcal{H}=\mathbb{C}^{d}$ is a function $R$ from $G$ to the group of unitary matrices on $\mathcal{H}$ such that $R(g)R(h)=R(gh)$. Maschke’s theorem asserts that for any $G$-representation $R$ there exists a unitary $U$ that simulaneously block-diagonalizes all $R(g)$ as:
$UR(g)U^{\dagger}=\bigoplus_{\alpha}I_{{\rm\mathbb{mult}}(r^{\alpha}_{G},R)}\otimes r^{\alpha}_{G}(g),\,$ (1)
[…]
  1. 1.
    Kostka Numbers: the multiplicity of the trivial $S_{\mu}$-irrep in the restriction of an $S_{n}$-irrep:
    $K^{\mu}_{\nu}={\rm\mathbb{mult}}(r^{\rm triv}_{S_{\mu}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{\mu}}),$ (2)
    where $\nu,\mu\vdash n$ and $S_{\mu}=\bigtimes_{i}S_{\mu_{i}}$ is a Young subgroup. $K^{\mu}_{\nu}$ also counts the number of semistandard Young tableaux of shape $\nu$ and content $\mu$. They play role in quantum tomography (see for example Ref. [6]).
  2. 2.
    Littlewood-Richardson Coefficients: the multiplicty of $S_{a}\times S_{b}$-irreps in an $S_{n}$-irrep for $a+b=n$:
    $c^{\nu}_{\lambda\mu}$ $={\rm\mathbb{mult}}(r^{\lambda}_{S_{a}}\otimes r^{\mu}_{S_{b}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{a}\times S_{b}}).$ (3)
    By Schur-Weyl duality [7, 8] this is also a multiplicty of $\textsf{SU}(d)$-irreps in the tensor product of two other $\textsf{SU}(d)$-irreps. The coefficient plays a role in angular momentum recoupling theory (for $d=2$, see for example Refs. [9, 10, 11]) or nuclear and particle physics (for example for $d=3$, the “eight-fold” way [12]).
  3. 3.
    Plethysm Coefficients: the multiplicity of irreps of the wreath product subgroup $S_{c}\wr S_{d}\subset S_{n}$ for $cd=n$ in an $S_{n}$-irrep:
    $a^{\nu}_{\lambda\mu}$ $={\rm\mathbb{mult}}(r^{\lambda\mu}_{S_{c}\wr S_{d}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{c}\wr S_{d}})\,,$ (4)
    where $\lambda,\mu,\nu\vdash c,d,n$. It is also the multiplicity of $\textsf{SU}(d)$-irreps in the composition of a pair of them. They appear in the quantum marginal problem for distinguishable particles [13, Sec. 7].
Our construction also subsumes the algorithm for Kronecker coefficients which previously appeared in Ref. [1]. Kronecker coefficients are multiplicities of $S_{n}$-irreps in the tensor products of $S_{n}$-irreps:
$g_{\lambda\mu\nu}$ $={\rm\mathbb{mult}}(r^{\nu}_{S_{n}},r^{\lambda}_{S_{n}}\otimes r^{\mu}_{S_{n}}),$ (5)
and play role in the quantum marginal problem [8, 14].
Computation of Plethysm and Kronecker coefficients is worst-case #$\#{\mathsf{P}}$-hard if we treat $n$ in $S_{n}$ as the input size (this corresponds to the unary input encoding) [15, 16] and it has been conjectured that the same is true for the Littlewood-Richardson and Kostka coefficients [17, Conjecture 5.11]  [^1] The computation of Littlewood-Richardson coefficients (to which the Kostka number computation reduces parsimonously) is #$\#{\mathsf{P}}$-hard for binary encoding by a reduction to Knapsack [47]. Binary encoding encodes the length of each row in the input integer partition in binary; this means that for a fixed input size $n$, the order of the underlying symmetric group ranges from $S_{n}$ to $S_{2^{n}}$. Fixing the order of $S_{n}$, as we do here, is equivalent to using unary encoding of the input. It is open if Kostka and LR-coefficients are #$\#{\mathsf{P}}$-hard with unary encoding. See [17].. While it would be surprising to find a quantum algorithm that computes the coefficients efficiently on all inputs, we give quantum algorithms that are efficient on some inputs and analyze their performance.
Quantum Algorithms.–
Let $G$ be a finite group and $R$ be a $G$-representation on $\mathbb{C}[G]$ with orthonormal basis {|$\{\ket{g}\}_{g\in G}$ acting as:
$R(h)\ket{g}$ =|$=\ket{hg},\quad\text{for all $h,g\in G$.}$ (6)
Such representation is called left-regular representation. A transformation $U$ that maximally block-diagonalizes this representation as in Eq. (1) is the Quantum Fourier Transform (QFT); we write $G$-QFT whenever we need to make the dependence on $G$ explicit. It is defined as:
$\langle\alpha,i,j|U|g\rangle$ $=\sqrt{\frac{d_{\alpha}}{|G|}}\innerproduct{i}{r_{G}^{\alpha}(g)|j},$ (7)
where $r_{G}^{\alpha}(g)$ is a $G$-irrep labelled by $\alpha$ and $\innerproduct{i}{r_{G}^{\alpha}(g)|j}$ for $i,j\in[d_{\alpha}]$ is its $i,j$-th matrix element. Efficient quantum algorithms for $G$-QFT are known for $G=S_{n}$ [19] and some other finite groups  [20]. We use QFT over the Young subgroup $S_{\mu}=\bigtimes_{i}S_{\mu_{i}}$ and the wreath product group $S_{a}\wr S_{b}$ from Ref. [20] (see Supp. Mat.).
Our algorithm also relies on the efficient preparation of states of the form:
$\rho^{\alpha}_{G}$ $=U^{\dagger}\left(\ket{\alpha}\bra{\alpha}\otimes\frac{I_{d_{\alpha}}}{d_{\alpha}}\otimes\frac{I_{d_{\alpha}}}{d_{\alpha}}\right)U=\frac{\Pi^{G}_{\alpha}}{d_{\alpha}^{2}},$ (8)
where $I_{d_{\alpha}}$ is the $d_{\alpha}\times d_{\alpha}$ identity matrix and $\Pi^{G}_{\alpha}$ is the projector onto the maximally isotypic subspace $\alpha$ in the regular representation of $G$; where the maximally isotypic subspace of a given type is the subspace of $\mathcal{H}$ containing all irreps of such type. Eq. (8) shows that an efficient $G$-QFT is sufficient, but possibly not necessary, to prepare this state. Equipped with this, we prove:
Theorem 1.
Let $H\subseteq G$. Let $\mathcal{H}$ be the size of a quantum circuit for $H$-QFT, $\mathcal{R}$ be the size of the circuit for preparing $\rho^{\alpha}_{G}$ in Eq. (8) and $\mathcal{D}$ be the size of a circuit that implements a controlled action of $G$-regular representaton. Then there is an $\mathcal{O}((\mathcal{H}+\mathcal{R}+\mathcal{D})(d_{\alpha}/d_{\beta})^{2})$ quantum algorithm for the following problem: given a pair of irrep labels $(\alpha,\beta)$ as an input, compute (with constant probability) ${\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow_{H}^{G})$ in time  [^2] The irreps $\alpha,\beta$ can be labeled with $O(\log|G|)$-many bits each, so that the input size $n$ relates nontrivially to the order of the group $G$. Since there are at most $|G|$ distinct irreps, such labelling always exists..
Proof.
The algorithm is a variation on Harrow’s generalized phase estimation [22] shown in Fig. 1. Let $G$ be a finite group and let $H\subseteq G$. The algorithm uses two registers: a target register $\mathcal{T}$ and a control register $\mathcal{C}$. It inputs $\rho^{\alpha}_{G}$ from Eq. (8) to the target register and a uniform superposition over $H\subseteq G$ to the control, prepared by applying an inverse $H$-QFT to the state |$\ket{\text{triv},1,1}$, where triv labels the trivial irrep of $H$. The algorithm then applies a controlled action of the $G$-regular representation $R$ to $\rho^{\alpha}_{G}$. Since $R$ is controlled by a register that contains a superposition over the subgroup $H$, this acts as a restricted representation $R\downarrow_{H}^{G}$ and decomposes into irreducible representations of $H$. The algorithm concludes by measuring the $H$-irrep label on the control register. All of this can be described by:
$E_{\beta}=\frac{1}{\sqrt{|H|}}\sum_{h\in H}\Bigg{[}(\outerproduct{\beta}{\beta}\otimes I_{d_{\beta}}\otimes I_{d_{\beta}})V\ket{h}\Bigg{]}_{\mathcal{C}}\otimes\Bigg{[}R(h)\Bigg{]}_{\mathcal{T}},$ (9)
where $V$ is the $H$-QFT and $\beta$ labels inequivalent irreps of $H$. A measurement outcome $\beta$ appears with probability:
$p_{\alpha}(\beta)$ $=\text{Tr}\left[E_{\beta}\rho^{\alpha}_{G}E_{\beta}^{\dagger}\right]={\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow^{G}_{H})\frac{d_{\beta}}{d_{\alpha}}.$ (10)
Because ${\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow^{G}_{H})$ is an integer, we can get its exact value in $\mathcal{O}((d_{\alpha}/d_{\beta})^{2})$-many shots with high probability by the Chernoff-Hoeffding bound (see Supp. Mat.). The algorithm computes ${\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow^{G}_{H})$ in $\mathcal{O}((\mathcal{H}+\mathcal{R}+\mathcal{D}))(d_{\alpha}/d_{\beta})^{2})$ time. ∎
Tab. 1 summarizes different choices of $H$ and $G$ where the quantum algorithm in Thm. 1 computes the coefficients in Eqs. (2) to (5). Note that nearly all cases considered have the same ambient group $G=S_{n}$. Since $\log|S_{n}|=\log n!\approx n\log n\in\widetilde{\mathcal{O}}(n)$, we parametrize the complexity of the algorithms below by $n$. We also parametrize the complexity of the problem in terms of dimensions of $S_{n}$-irreducible representations which can be computed in polynomial time in $n$ using the hook length formula. Following results Ref. [23], we assume that the $S_{n}$-QFT can be implemented in gate complexity $\mathcal{O}(n^{4})$.
Figure 1: The quantum algorithm in Thm. 1. The algorithm initializes the state from Eq. (8), applies a $G$-regular representation controlled by a uniform superposition over $H$ and weak-Fourier samples the control register.
Kostka Numbers.—
The Kostka number $K_{\nu}^{\mu}$ is the multiplicity of the trivial irreducible representation in the restriction of an $S_{n}$-irrep $r^{\nu}_{S_{n}}$ to $S_{\mu}$, $K^{\mu}_{\nu}={\rm\mathbb{mult}}(r_{S_{\mu}}^{{\rm triv}},r^{\nu}_{S_{n}}\downarrow_{S_{\mu}}^{S_{n}})$. A quantum algorithm for the Kostka numbers fixes $H=S_{\mu}\subseteq G=S_{n}$ for an integer partition $\mu\vdash n$, where $S_{\mu}$ is the Young’s subgroup $S_{\mu}:=\bigtimes_{i}S_{\mu_{i}}$. An efficient $S_{\mu}$-QFT can be setup from a tensor product of $S_{\mu_{i}}$-QFTs for each factor in $\mu\vdash n$ and its complexity will be comparable to $S_{n}$-QFT. Theorem 1 then gives the label of the trivial irreducible representation of $S_{\mu}$ with probability (c.f. Eq. (10)):
$p_{\nu,\mu}({\rm triv})$ $=K^{\mu}_{\nu}/d_{\nu}.$ (11)
This suffices to compute the Kostka number in $\mathcal{O}(n^{4}d_{\nu}^{2})$ Computing Kostka numbers $K^{\mu}_{\nu}$ on an input $\mu,\nu$ in unary is conjectured to be #$\#{\mathsf{P}}$-complete in [17, Conjecture 5.11] and one could hope that restricting $d_{\nu}\leq{\rm poly}(n)$ remains classicaly hard. We show that this is not the case:
Theorem 2.
There is a deterministic algorithm that on an input $\mu,\nu\vdash n$ in unary, such that $d_{\nu}\leq{\rm poly}(n)$, computes the Kostka number $K_{\nu}^{\mu}$ in polynomial time in $n$.
See the Supp. Mat. for a proof.
Littlewood-Richardson Coefficients.–
The quantum algorithm for Littlewood-Richardson (LR) coefficients fixes $H=S_{a}\times S_{b}\subseteq G=S_{n}$ for $a+b=n$. The irreducible representations of $S_{a}\times S_{b}$ are a tensor product of irreducibles $r_{S_{a}}^{\lambda}\otimes r_{S_{b}}^{\mu}$, labelled by pairs $\alpha=(\lambda,\mu)$ with $\lambda,\mu\vdash a,b$. We have $c^{\nu}_{\lambda\mu}={\rm\mathbb{mult}}(r^{\lambda}_{S_{a}}\otimes r^{\mu}_{S_{b}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{a}\times S_{b}})$, and from Thm. 1, the algorithm samples from:
$p_{\nu}(\lambda,\mu)$ $=c^{\nu}_{\lambda\mu}\frac{d_{\lambda}d_{\mu}}{d_{\nu}}.$ (12)
This computes the LR coefficient in $\mathcal{O}\left(n^{4}\left(\frac{d_{\nu}}{d_{\lambda}d_{\mu}}\right)^{2}\right)$.
LR coefficients count combinatorial objects called the LR tableaux. A skew shape $\nu/\lambda$ is a diagram where $\lambda$ is substracted from $\nu$ and a skew semistandard Young Tableau is a skew shape filled with integers such that the entries are non-decreasing along every row and strictly increasing along every column. The LR tableau is a skew Young tableau of shape $\nu/\lambda$ whose row-reading evaluates to $\mu$. The LR coefficient $c^{\nu}_{\lambda\mu}$ is the number of such tableaux; this fact is called the LR rule. See [24, 25, 26] and the Supp. Mat. for details (esp. definition of ‘row reading’). The quantum algorithm in Thm. 1 is only efficient for $c^{\nu}_{\lambda\mu}\leq{\rm poly}(n)$ in which case it counts only up to ${\rm poly}(n)$ many LR tableaux. In analogy to Thm. 2, we expect that there is a classical algorithm that efficiently computes LR coefficients under this constraint:
Hypothesis 1.
There is a polynomial-time randomized algorithm for computing LR coefficients $c^{\nu}_{\lambda\mu}$ for $\nu\vdash n$, $\lambda\vdash a$ and $\mu\vdash b$ for $a+b=n$ and $d_{\nu}/d_{\lambda}d_{\mu}\leq{\rm poly}(n)$.
See also [27]. If true, this rules out superpolynomial speedups for computing LR coefficients by Thm. 1.  [^3] We make a distinction between a hypothesis and a conjecture. By a hypothesis, we mean claim that we strongly believe in but cannot prove. A conjecture labels a claim that we cannot prove and have no any strong beliefs about.
Plethysm Coefficients.— The algorithm for Plethysm coefficients sets $H=S_{c}\wr S_{d}\subseteq G=S_{n}$ for $cd=n$. An efficient implementation of the QFT over this group is known only for $|S_{c}|\leq{\rm poly}(n)$ [20]. Under this restriction, the quantum algorithm in Thm. 1 samples efficiently from a distribution:
$p_{\nu}(\lambda,\mu)$ $=a^{\nu}_{\lambda\mu}\frac{d_{\lambda}d_{\mu}^{d}}{d_{\nu}},$ (13)
where $a^{\nu}_{\lambda\mu}={\rm\mathbb{mult}}(r^{\lambda,\mu}_{S_{c}\wr S_{d}},r^{\nu}_{S_{n}}\downarrow^{S_{n}}_{S_{c}\wr S_{d}})$ is the Plethysm coefficient. It is not known if Plethysm coefficients count a well defined set of combinatorial objects (Prob. 9 on Stanley’s list [29]), and the simulation strategy of Thm. 2 doesn’t straightforwardly apply [30, 16, 31]. The classical complexity of computing the Plethysm coefficients was recently studied in Ref. [16] where it was shown that the problem is in ${\mathsf{P}}$ if the number of parts $\ell(\lambda)$ is constant and $\mu={\rm triv}$, but becomes #$\#{\mathsf{P}}$-hard even for fixed $d=3$ and $\ell(\lambda)$ unrestricted. While this suggests that the complexity of the problem is controlled by the parameter $c$, its complexity for $|S_{c}|={\rm poly}(n)$ ($c\sim\log(n)$) remains, to our best knowledge, open. We thus state:
Conjecture 1 (See also [27]).
There is no polynomial time randomized algorithm that (with high probability) computes $a^{\nu}_{\lambda\mu}$ for $\lambda\vdash c$ and $\mu\vdash d$, such that $cd=n$, $c\leq{\rm poly}(\log(n))$ and $d_{\lambda}d_{\mu}^{d}/d_{\nu}\geq 1/{\rm poly}(n)$.
If true, the quantum algorithm for Plethysm coefficients gives a superpolynomial quantum speedup. The fact that an analog of Thm. 2 does not apply here supports this, but the constraints on the conjectured computationally hard regime for the problem are very stringent and can make the restricted computational problem significantly easier. We have little intuition whether the conjecture is true or false, besides the recent obstructions proved in [27] as we were finishing this manuscript.
Kronecker Coefficients.— Given $\lambda,\mu,\nu\vdash n$, the Kronecker coefficient $g_{\lambda\mu\nu}$ is defined as $g_{\lambda\mu\nu}={\rm\mathbb{mult}}(r^{\lambda}_{S_{n}}\otimes r^{\mu}_{S_{n}},r^{\nu}_{S_{n}})$. The algorithm in Ref. [32] is a special case of Thm. 1, for the choice $H=S_{n}\subseteq G=S_{n}\times S_{n}$. It samples from the output distribution:
$p_{\lambda\mu}(\lambda)$ $=\frac{d_{\nu}}{d_{\lambda}d_{\mu}}g_{\lambda\mu\nu}\,,$ (14)
and computes $g_{\lambda\mu\nu}$ (with high constant probability) in time $\mathcal{O}\left(n^{4}(d_{\lambda}d_{\mu}/d_{\nu})^{2}\right)$. Similarly to Plethysms, it remains a major open question if the Kronecker coefficients count some simple set of combinatorial objects (Stanley’s Prob. 10 [29]), which again precludes straightforward application of the simulation strategy of Thm. 2. The classical complexity of computing the Kronecker coefficients was studied in Ref. [31] which identified the extensive number of parts as a potential obstruction to their efficient computation. They show that $g_{\lambda\mu\nu}$ can be computed in time:
$\mathcal{O}\Big{(}L\log(N)\Big{)}+\Big{(}L\log(M)\Big{)}^{\mathcal{O}(L^{3}\log L)},$ (15)
where $L=\max\Big{(}\ell(\lambda),\ell(\mu),\ell(\nu)\Big{)}$, $N=\max(\lambda_{1},\mu_{1},\nu_{1})$ is the largest length of the first row and $M=\max(\lambda_{2},\mu_{2},\nu_{2})$ is the largest length of the second row among these partitions [31, Thm. 3.5]. Even for fixed $L$, the algorithms of Refs. [33, 34] compute the Kronecker coefficient in polynomial time but become inefficient when the length of the input partitions varies with input size. The algorithm in Ref. [35] computes Kronecker coefficient by reduction to vector partition function problem. All of these algorithms are based on Barvinok’s algorithm for counting the lattice points in convex polytopes (Refs. [36, 37, 38, 39, 40, 41, 42]).
The quantum algorithm in Thm. 1 is for all practical purposes outperformed by such algorithms on some inputs. For example, Ref. [35, Ex. 3.4] computes $g_{\lambda\mu\nu}=391$ for $\lambda=[140^{2}],\;\mu=[70^{4}],\;\nu=[57^{3},33^{3},10]$ for which $d_{\nu}/(d_{\lambda}d_{\mu})\approx 1.8\times 10^{-31}$ on which the quantum algorithm in Thm 1 requires an order of $10^{31}$ samples. On the other hand, it runs in polynomial when $d_{\mu}\leq{\rm poly}(n)$ and $\lambda=\nu$ is arbitrary. We give an explicit construction of a family of such inputs in the Supp. Mat..
We originally took the length obstruction in Eq. (15) as a support for the conjecture of Ref. [1] that the algorithm for the Kronecker coefficients can outperform classical algorithms for triples of partitions with large number of parts. This lead us to state:
Conjecture 2 (Disproved in [27]).
There is no polynomial time randomized algorithm that (with high probability) computes $g_{\lambda\mu\nu}$ for $\mu,\nu,\lambda\vdash n$ and $d_{\mu}\leq{\rm poly}(n)$.
However, Greta Panova disproved this conjecture by extending the result of Ref. [33] by showing that whenever $d_{\mu}\leq n^{k}$, there is a $\tilde{\mathcal{O}}(n^{4k^{2}+1})$ classical algorithm for the problem [27, Theorem 1.]. A more general form of this conjecture (as discussed in [1]) narrowly escapes this disproof but requires fine-tuned inputs, so we take Panova’s result as a strong evidence against superpolynomial quantum speedups for this problem (see footnote  [^4] There is no polynomial time randomized algorithm that (with high probability) computes $g_{\lambda\mu\nu}$ for $\mu,\nu,\lambda\vdash n$ and $d_{\nu}/(d_{\lambda}d_{\mu})\geq 1/{\rm poly}(n)$. The key challenge then becomes to find a nicely parametrized set of inputs such that $d_{\nu},d_{\lambda},d_{\mu}$ are all superpolynomial, their ratio is bounded by a polynomial and the partitions do not fall into one of the special cases that can be simulated easily (such as hooks). As long as $\lambda,\mu$ have number of parts that scale with $n$, the quantum algorithm can do better than the bound in Eq. (15) as well as the state of the art classical algorithms, but the new result in [27] supersedes this obstruction). Interestingly, the gate complexity of our algorithm is ${\mathcal{O}}(n^{4}(d_{\mu}d_{\lambda}/d_{\nu})^{2})$ so that a ${\mathcal{O}}(n^{4+2k})$ vs. $\tilde{\Omega}(n^{4k^{2}+1})$ polynomial speedup for $d_{\lambda}=d_{\nu}$ and $d_{\mu}\leq n^{k}$ is possible, assuming that the scaling in [27, Theorem 1.] is tight. As this asks for additional scrutiny, we pose:
Question 1.
Is there a classical algorithm that computes $g_{\mu\nu\lambda}$ in time ${\mathcal{O}}(n^{2k+4})$ whenever $d_{\mu}\leq n^{k}$?
See the Supp. Mat. for an explicit construction of a nontrivial family of inputs on which our algorithm is efficient, additional analysis, discussion and numerical study.
Induction Algorithm.–
Frobenius’s reciprocity [7]:
${\rm\mathbb{mult}}(r^{\beta}_{H},r^{\alpha}_{G}\downarrow_{H}^{G})={\rm\mathbb{mult}}(r^{\alpha}_{G},r^{\beta}_{H}\uparrow_{H}^{G})\,,$ (16)
suggests an alternative to Thm. 1 that computes multiplicities by Fourier sampling over induced representations:
Theorem 3.
Let $H\subseteq G$. If there is a quantum circuits for $G$-QFT with $\mathcal{G}$ many gate, a quantum circuit with $\mathcal{R}$ many gates for preparing $\rho^{\beta}_{H}$ in Eq. (8), then there is a $\mathcal{O}((\mathcal{G}+\mathcal{R})(d_{\beta}|G|/d_{\alpha}|H|)^{2})$ quantum algorithm that, given a pair of irrep labels $(\alpha,\beta)$ as an input, computes (with high probability) ${\rm\mathbb{mult}}(r^{\alpha}_{G},r^{\beta}_{H}\uparrow_{H}^{G})$ in time .
Proof.
See Fig. 2. The algorithm uses two registers, an $H$-regular rep register $\mathcal{S}=\mathbb{C}[H]$ and a coset register $\mathcal{V}=\mathbb{C}[G/H]$. It prepares $\rho^{\beta}_{H}$ from Eq. (8) in $\mathcal{S}$ and $|H|\mathds{I}/|G|$ in $\mathcal{V}$ and applies an ‘embedding’ unitary $U_{E}$ to $\mathcal{S}\otimes\mathcal{V}$: $U_{E}\Big{(}\ket{t}_{\mathcal{V}}\otimes\ket{h}_{\mathcal{S}}\Big{)}=\ket{th},$ where $t$ is a representative of the coset of $H$ in $G$ and $th$ is an element of $G$. This unitary can be implemented efficiently as long as the multiplication of the group elements in $G$ is efficient an we assume that it’s complexity is negligible compared to both $\mathcal{G}$ and $\mathcal{R}$. The algorithm applies a $G$-QFT to $\mathcal{S}\otimes\mathcal{V}$ and measures the irrep label. The outcome $\alpha$ appears with probability:
$q_{\beta}(\alpha)$ $=\frac{|H|d_{\alpha}}{|G|d_{\beta}}{\rm\mathbb{mult}}(r^{\alpha}_{G},r^{\beta}_{H}\uparrow_{H}^{G}).$ (17)
Following the analysis in Thm. 1, we can find ${\rm\mathbb{mult}}(r^{\alpha}_{G},r^{\beta}_{H}\uparrow_{H}^{G})$ in $\mathcal{O}((\mathcal{G}+\mathcal{R})(d_{\beta}|G|/d_{\alpha}|H|)^{2})$ with high probability. See Supp. Mat. for details. ∎
Figure 2: The quantum algorithm in Thm. 3. The algorithm applies the embedding unitary $U_{E}$ to the subgroup and coset register and weak Fourier samples.
The sampling cost of the restriction algorithm in Thm. 1 is $d_{\alpha}/d_{\beta}$, compared to $(d_{\beta}|G|)/(d_{\alpha}|H|)$ in Thm. 3 and the restriction is preferable whenever $(d_{\alpha}/d_{\beta})^{2}<|G|/|H|$. For $\beta={\rm triv}$, Eq. (17) is the distribution of Weak Fourier sampling from the hidden subgroup state, appearing in the context of the Hidden Subgroup Problem (HSP). The computational task in HSP is of course different from computing the multiplicity, but the connection suggests generalizations of HSP to $\beta\neq{\rm triv}$ we explore in future work.
Discussion.–
We gave quantum algorithms for Kostka, Littlewood-Richardson, Kronecker and Plethysm coefficients, efficient on a nontrivial subset of inputs. We gave a polynomial-time classical algorithm for computing of the Kostka numbers $K^{\mu}_{\nu}$ whenever $d_{\nu}\leq{\rm poly}(n)$, ruling out a superpolynomial quantum speedup for this problem and conjectured the existence of a similar classical algorithm for the Littlewood-Richardson coefficients. We explained why such classical algorithm does not straightforwardly extend to the Plethysm and Kronecker coefficients and conjectured that our quantum algorithms efficiently solve problems that are hard for classical algorithms. Going against our thesis is a result in Ref. [44] in which Jordan presented a quantum algorithm for approximating $S_{n}$ characters to a similar precision as our Thm. 1 and gave an efficient classical algorithm based on the Roichman’s rule [45]. Our Conjecture 2 followed a similar fate and was disproved by an algorithm in [27, Theorem 1.]. The runtime of this algorithm is $\tilde{\mathcal{O}}(n^{4k^{2}+1})$ which, if tight, suggests a possible ${\mathcal{O}}(n^{4+2k})$ vs $\tilde{\Omega}(n^{4k^{2}+1})$ speedup on inputs with $d_{\nu}=d_{\lambda}$ and $d_{\mu}\leq n^{k}$. It is thus important to address Question 1 and understand if the large polynomial gap for computing Kronecker coefficient closes. Conjecture 1 was put under similar scrutiny: there are significant limitations to regimes where the algorithm can offer a superpolynomial speedup, but large polynomial speedups may be still possible. We note that, Ref. [46] showed that some of the multiplicity coefficients considered here appear naturally in spin chains. It remains an interesting question if this approach could provide more insight into this class of quantum algorithms.
Acknowledgments.–
We would like to thank David Gosset, Sergey Bravyi, Louis Schatzki, Greta Panova, Marco Cerezo, Marcos Crichigno, Samuel Slezak, Diego Garcia-Martín, Frederic Sauvage and Dmitry Grinko for discussions. Last but not least, we would like to thank an anonymous QIP referee who helped us to significantly improve the presentation of this paper, especially regarding the presentation of the correctness of the $S_{n}$-QFT - the previous version contained many mistakes. We had decided to rewrite that appendix altogether and if you ever get a chance to read this, we are seriously grateful for the detailed feedback and sorry for not updating our manuscript earlier. M.L. was supported by the Center for Nonlinear Studies at LANL and by the Laboratory Directed Research and Development (LDRD) program of LANL under project number 20230049DR.